Please show me how to slove problems
(5,-2), m=3
(5,4), m=-5
(-3,-1),m= -2/3

Answer by checkley71(8403)

Y=mX+b
-2=3*5+b
-2-15=b
b=-17 THUS THE EQUATION IS
Y=3X-17
---------------------------------------
4=-5*5+b
4+25=b
b=29 THUS THIS EQUATION IS
Y=-5X+29
----------------------------------------
-1=-2/3*-3+b
-1-6/3=b
-1-2=b
b=-3 THUS THE EQUATION IS
Y=-2/3X-3